OpenStudy (anonymous):

How many moles of ethane (C2H6) would be needed to react with 62.3 grams of oxygen gas? Unbalanced equation: C2H6 + O2 → CO2 + H2O

OpenStudy (anonymous):

balenced equation is C2H6+5O2 yield 2CO2+3H2O

OpenStudy (anonymous):

the balanced equation is 2(C2H6) + 7(O2) >>>> 4(CO2)+6(H2O)

OpenStudy (anonymous):

If you do the stoichiometry you get an answer of 0.556 mol of ethane (C2H6)

OpenStudy (anonymous):

(62.3gO2) * (1mol O2 / 31.998 g O2) * (2mol C2H6 / 7mol O2) = 0.566 mol ethane

OpenStudy (anonymous):

with sig figs --> .389 mol C2H6 \[62.3g \space O_{2} \times 1mol O_{2} \times 4molC_{2} H_{6} \over 32.00g \space O_{2} \times 7mol \space O_{2}\]

OpenStudy (anonymous):

@tktechie where did you get the 4 mol of C2H6 in your stoichiometry?

OpenStudy (anonymous):

@ SalvatoreTRM Oh crud! That was a typo.. and the answer i gave was wrong, too. the answer was based off of the original balace that the asker gave. I didn't think to check it.